### v1

# Lifshitz tail states in non-Hermitian disordered photonic lattices

Abstract
In lattices with uncorrelated on-site potential disorder, Anderson localization near the band edges can exhibit anomalously weak localization in the form of Lifshitz tail states. These states correspond to clusters of contiguous sites with nearly identical on-site energies, allowing excitations to extend significantly beyond the characteristic localization length determined by the inverse of Lyapunov exponent. Since Lifshitz tail states are rare events, with an exponentially small density of states, they are typically considered of limited practical importance. In this work, we demonstrate that when Anderson localization is induced by disorder in an imaginary on-site potential, Lifshitz tail states can dominate the system’s dynamics and become experimentally observable. This phenomenon is illustrated through the Anderson-Bernoulli model in a non-Hermitian photonic lattice, shedding light on the unique interplay between disorder and non-Hermiticity in such systems.
journal: opticajournal
Introduction. Disorder-induced phenomena in wave systems, particularly Anderson localization [1, 2, 3, 4, 5], have attracted substantial interest across various fields, including condensed matter physics, optics, and acoustics. Anderson localization occurs when disorder is strong enough to disrupt wave transport, leading to the spatial confinement of eigenstates. This phenomenon has been observed in numerous physical systems, with photonics serving as a particularly captivating laboratory for investigating and observing such effects [6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. In one-dimensional lattices with uncorrelated on-site potential disorder, all eigenstates are exponentially localized, characterized by a localization length which is expected to be inversely related to the Lyapunov exponent. However, near the spectral band edges, certain eigenstates exhibit anomalously weak localization, forming the so-called Lifshitz tail states [4, 16, 17, 18, 19, 20]. These rare states emerge from clusters of contiguous sites possessing nearly identical on-site energies, which facilitate significantly extended excitation compared to the typical localization length [4, 19]. This kind of "delocalization" at the tail can be detected by a marked deviation of the inverse participation ratio of the wave function from its Lyapunov exponent near the band edges [19], a phenomenon very distinct from the well-established mobility edge (localization-delocalization transition). Since Lifshitz tail states are extremely rare events and correspond to exponentially suppressed spectral densities [4], they appear of limited relevance in conventional Hermitian systems and have largely been relegated to theoretical discussions [18, 19, 20].
In recent years, non-Hermitian systems –characterized by complex potentials– have emerged as a fertile ground for novel wave phenomena, driven by their ability to model energy gain, loss, and nonreciprocal effects [21]. Non-Hermitian extensions of Anderson localization have uncovered phenomena distinct from their Hermitian counterparts, including the non-Hermitian delocalization transition [22, 23, 24, 25] and a form of universal non-Hermitian transport [26, 27, 28, 29, 30]. In these systems, wave propagation can occur in a jumpy manner even when the eigenstates of the Hamiltonian are exponentially localized [26, 30]. Lifshitz tail states are also present in non-Hermitian disordered lattices [31, 32]. Although their physical significance and practical relevance have remained uncertain [32], a recent study [30] highlighted the role of the density of state tails in establishing the dynamical scaling in non-Hermitian transport.
In this Letter, we show that Lifshitz tail states, despite being extreme events, can dominate the dynamical behavior of disordered non-Hermitian lattices of finite size. Specifically, we focus on Anderson localization induced by imaginary on-site potential Bernoulli disorder in non-Hermitian photonic lattices. We demonstrate that, in such systems, Lifshitz tail states not only become more prominent but also play a pivotal role in shaping the overall system behavior. These findings challenge conventional views of Lifshitz tail states as negligible and highlight their experimental observability in non-Hermitian settings. To illustrate this phenomenon, we analyze the Anderson-Bernoulli model, a paradigmatic framework for studying localization in a two-component alloy [18, 20, 33, 34], adapted here to a non-Hermitian photonic lattice. Our results provide insights into the dynamics of non-Hermitian systems with disorder and reveal potential avenues for harnessing non-Hermitian models to visualize rare events like Lifshitz tail states.
Model and Lifshitz tail states. We consider a one-dimensional tight-binding lattice of size $L$ with a random on-site potential. The Schrödinger equation for a single particle hopping on the lattice reads
$i\frac{d\psi_{n}}{dt}=J(\psi_{n+1}+\psi_{n})+V_{n}\psi_{n}=\sum_{m=1}^{L}\mathcal{H}_{n,m}\psi_{m}$ (1)
where $\psi_{n}(t)$ is the wave amplitude at the $n$-th lattice site ($n=1,2,..,L$), $J$ is the hopping rate, $V_{n}$ is the on-site random potential, and $\psi_{0}(t)=\psi_{L+1}(t)=0$ for open boundary conditions. $V_{n}$ are assumed to be independent random variables with the same probability density distribution. For the Bernoulli-Anderson model [20, 33, 34], $V_{n}$ can take only two values, $V_{n}=W$ with probability $p$ and $V_{n}=0$ with probability $(1-p)$. Hence the on-site potential is a stochastic sequence of 0 and $W$. In the non-Hermitian Anderson-Bernoulli model, the potential is imaginary, namely $W=-i\gamma$ where $\gamma>0$ is the particle loss rate at site $n$. In the photonic context, this model basically describes a coherent random walk of a photon on a lattice with traps [35], which can be experimentally implemented in synthetic temporal mesh lattices (see e.g. [15, 26, 36, 37] and reference therein).
The eigenvalues $\lambda$ of the matrix Hamiltonian $\mathcal{H}$ are complex with ${\rm Im}(\lambda)\leq 0$, and we are interested in the behavior of $\lambda$ and corresponding eigenvectors $u_{n}$ in the large system size ($L\rightarrow\infty$) limit. For a large loss rate $\gamma\gg J$, the eigenvalues $\lambda$ and corresponding eigenvectors $u_{n}$ form two distinct groups, one with ${\rm Im}(\lambda)\simeq-\gamma$ corresponding to short-lived states, and the other one with ${\rm Im}(\lambda)\simeq 0^{-}$ corresponding to long-lived states; see Fig.1(a). The eigenvectors of the former group have their excitations mainly in the lossy sites, where $V_{n}=-i\gamma$, while the latter group of eigenstates have their excitations mostly in the lossless sites, where $V_{n}=0$. The energy spectrum of both groups is symmetric about the real axis ${\rm Re}(\lambda)=0$. In fact, it can be readily shown that, if $u_{n}$ is an eigenfunction of $\mathcal{H}$ with eigenvalue $\lambda$, then $(-1)^{n}u^{*}_{n}$ is an eigenvector of $\mathcal{H}$ with eigenvalue $-\lambda^{*}$.
The dynamical evolution of the system is determined by the eigenstates possessing the longest lifetime [26, 30], i.e. the eigenstates with ${\rm Im}(\lambda)$ closest to zero, which we call the Lifshitz tail eigenstates. They appear in pairs with the same imaginary part of the eigenvalue but opposite real part. As shown in the inset of Fig.1(a), the Lifshitz tail eigenstates with the longest lifetime are those with real part of energy $\lambda$ approaching the band edges ±$\pm 2J$ in the real part of the energy spectrum, and basically correspond to the low-energy eigenstates of the "particle-in-a-box" problem [20, 34], with the largest box length $l\sim L\rightarrow\infty$ as an extreme event. In fact, a realization of on-site potential $V_{n}$ can be viewed as random sequences of zeros (with stochastic length $l$) separated by sharp complex potential barriers $-i\gamma$ (the traps), i.e. stochastic sequences of potential wells of random length $l$ separated by infinite barriers in the $\gamma/J\gg 1$ limit. The Lifshitz tail eigenstates have their excitations mostly confined in the potential wells of large size $l$, for which the eigenstate $u_{n}$ can be approximated by the ground state of the infinite potential well, $u_{n}\simeq\sqrt{2/(l+1)}\sin[n\pi/(l+1)]$[Fig.1(b)]. The corresponding eigenvalue $\lambda$ is close to the ground-state energy of the potential well and can be estimated from a perturbative analysis as (see the Supplemental document for technical details)
Figure 1: (a) Behavior of the energy spectrum $\lambda$ of the matrix Hamiltonian $\mathcal{H}$ in complex plane for a given realization of the random Bernoulli potential $V_{n}$. Parameter values are $\gamma/J=5$, $p=0.2$ and $L=5000$. The spectrum is symmetry around the imaginary axis ${\rm{Re}}(\lambda)=0$ and can be grouped in two sets: the eigenvalues with small imaginary parts ($|{\rm Im}(\lambda)|\sim J^{2}/\gamma$), corresponding to short-lived eigenstates, and the eigenvalues with large imaginary parts ($|{\rm Im}(\lambda)|\sim\gamma$), corresponding to long-lived eigenstates. The Lifshitz tail eigenstates are those with eigenvalues $\lambda$ close to the upper left and right corners, possessing extremely long lifetimes. The insets in (a) shows an enlargement of the spectrum near the upper right corner. The dashed curve in the panel depicts the loci of the eigenenergies of the Lifshitz tail states predicted by the "particle-in-a-box" model [Eq.(2)]. (b) Numerically-computed behavior of the Lifshitz tail state possessing the longest lifetime (modulus of $|u_{n}|$, red circles). The position of the traps is also shown for comparison. This eigenstate is well approximated by the ground state of a potential well of size $l_{m}$ (dashed rectangular curve), corresponding to the longest sequence of consecutive zeros in the disorder realization $V_{n}$ (blue squares).
$\lambda=\lambda_{r}+i\lambda_{i}\simeq J\left(2-\frac{\pi^{2}}{l^{2}}\right)-i\left(\frac{1}{l^{3}}\frac{4\pi^{2}\gamma J^{2}}{\gamma^{2}+4J^{2}}\right).$ (2)
Since the probability to find in the sequence a potential box of size $l$ is $P_{l}=p(1-p)^{l}$, exponentially vanishing as $l\rightarrow\infty$, the Lifshitz tail states are rare, and in fact the densities of states $\rho(\lambda_{r})$ and $\rho(\lambda_{i})$ ($\lambda_{r,i}$ are the real and imaginary parts of $\lambda$) shows a typical exponential decay behavior at the band tails as in the Hermitian case [4, 20, 34]. A major feature of Lifshitz tail states is that their spatial extension, of order $\sim l$, is much larger and strongly deviates from the inverse of the Lyapunov exponent [19]: in fact, the latter remains finite at the band edges and describes the exponential decay tails of the wave function $u_{n}$ outside the potential well, while the potential box size $l$ can be as large as $\sim L$ in an extreme event. The "tail delocalization" effect of Lifshitz states can be captured by computing the inverse participation ratio (IPR) [19] of the wave function $u_{n}$ in a box of length $l$ , which reads
${\rm IPR}_{l}=\frac{\sum_{n=1}^{L}{|u_{n}|^{4}}}{\sum_{n=1}^{L}{|u_{n}|^{2}}}=\frac{3}{2}\left(\frac{1}{l+1}\right).$ (3)
In a finite system of size $L$, the box length $l$ is clearly limited by $L$. Indicating by $l=l_{m}$ the largest potential well size in the lattice for a given realization of $V_{n}$, clearly $l_{m}$ is a stochastic variable with some probability distribution function. The mean value $\bar{l}_{m}$ of $l_{m}$ can be computed using the extreme value theory of stochastic Bernoulli processes and takes the approximate form (see e.g. [38])
$\bar{l}_{m}\simeq\frac{\ln\left(\sigma pL\right)}{-\ln(1-p)}$ (4)
[…]
System dynamics and Lifshitz tail delocalization. The delocalization features of the Lifshitz tail states are hard to be observed in large-size Hermitian systems, since they are extremely rare states and, for a given realization of the disorder sequence $V_{n}$, their spatial locations in the lattice are unknown, rendering their excitation unfeasible. Conversely, in non-Hermitian lattices Lifshitz tail states not only become more prominent but also play a pivotal role in shaping the long-time system behavior. In fact, let us assume that at initial time all sites of the lattice are equally excited, i.e. let us assume $\psi_{n}(0)=(1/\sqrt{L})$. This condition ensures that Lifshitz tail states of the right edge band are surely excited, regardless of their unknown location and size in the lattice. The solution to the Schrödinger equation (1) reads
$\psi_{n}(t)=\sum_{\alpha=1}^{L}C_{\alpha}u_{n}^{(\alpha)}\exp(-i\lambda_{\alpha}t)$ (5)
where $\lambda_{\alpha}$ and $u_{n}^{(\alpha)}$ are the $L$ eigenvalues and corresponding eigenvectors of $\mathcal{H}$, and $C_{\alpha}$ are the spectral excitation amplitudes, given by $C_{\alpha}=\sum_{n=1}^{L}u_{n}^{(\alpha)}$, with $C_{\alpha}\simeq\sqrt{2/(1+l)}{\rm cotg}[\pi/(2l+2)]$ for a Lifshitz tail state of size $l$. Let us indicate by $\alpha=\alpha_{0}$ the index corresponding to the Lifshitz tail state with the longest lifetime, i.e. corresponding to the potential box of size $l=l_{m}$. Clearly, since all other states decay faster, in the long time limit $t\rightarrow\infty$ one has
$\psi_{n}(t)\simeq C_{\alpha_{0}}u_{n}^{(\alpha_{0})}\exp(-i\lambda_{\alpha_{0}}t)$ (6)
i.e. the long-time system’s behavior is dominated by the most rare Lifshitz tail state. The emergence of such a state from the system’s dynamics can be detected by monitoring the temporal behavior of the dynamical IPR, which is defined as
${\rm IPR}(t)=\frac{\sum_{n=1}^{L}|\psi_{n}(t)|^{4}}{\sum_{n=1}^{L}|\psi_{n}(t)|^{2}}.$ (7)
According to the previous analysis, the statistical average of ${\rm IPR}(t)$, i.e. averaged over a large number of realizations of the Bernoulli disorder, should converge to the value ${\rm IPR}_{\bar{l}_{m}}$ as given by Eqs.(3) and (4), i.e. one has
$\lim_{t\rightarrow\infty}\overline{{\rm IPR}(t)}={\rm IPR}_{\bar{l}_{m}}\simeq\frac{3}{2}\frac{-\ln(1-p)}{\ln(\sigma pL)}.$ (8)
Figure 2: (a) Numerically-computed behavior of the IPR of the Lifshitz tail state with the longest lifetime versus lattice size $L$ (circles); the IPR is averaged over 200 realizations of Bernoulli disorder. Parameter values are $p=0.2$ and $\gamma/J=5$. The solid curve depicts the theoretically predicted behavior based on Eqs.(3) and (4). (b) Dynamical evolution of $|\psi_{n}(t)|$ in a lattice of size $L=500$ for a single realization of Bernoulli disorder. Initial condition is $\psi_{n}(0)=1/\sqrt{L}$. The upper panel in (b) shows the behavior of $|\psi_{n}|$ at final propagation time $t/J=1200$ (red circles), and the realization of Bernoulli disorder (blue squares). The main peak corresponds to the ground state of the potential box with largest size $l=l_{m}$ of the disorder realization. (c) Numerically-computed behavior of the dynamical IPR (solid curve), averaged over 100 realizations of disorder. The dashed horizontal curve in the asymptotic theoretical value predicted by Eq.(8).
This theoretical prediction turns out to be in good agreement with the numerical results. As an example, Fig.2(b) shows the evolution of $|\psi_{n}(t)|$, normalized at each time step to its norm, in a lattice of size $L=500$ for a single realization of the Bernoulli disorder. At initial time, all sites of the lattice are equally excited. As one can clearly see, in the long-time dynamics the system approaches the Lifshitz tail state with the longest lifetime, which is centered at the site $n=314$ and of size $l_{m}\simeq 28$, as shown in the upper panel of Fig.2(b). Figure 2(c) shows the numerically-computed time evolution of the dynamical IPR, averaged over 100 different realizations of the Bernoulli disorder. Note that in the long time limit the mean value $\overline{{\rm IPR}(t)}$ asymptotically approaches a constant value which is in good agreement with the theoretically predicted value given by Eq.(8).
Lifshitz tail states in synthetic temporal photonic lattices. An experimentally-accessible platform for the visualization of Lifshitz tail states is provided by photonic random walks in a synthetic temporal lattice with traps (see e.g. [15, 26, 35, 36, 37]). The system consists of two fiber loops of slightly different lengths ±$\mathcal{L}\pm\Delta\mathcal{L}$, that are connected by a variable directional coupler with a coupling angle $\beta$. Amplitude modulators are placed in the two loops, which introduce stochastic losses to emulate a Bernoulli distribution of traps [35]. By discretizing the physical time as $t=t_{n}^{(m)}=mT+n\Delta T$, where $T=\mathcal{L}/c$ is mean travel time and $\Delta T=\Delta\mathcal{L}/c\ll T$ is the travel-time difference of light pulses in two loops, light dynamics is described by the set of discrete-time equations (see e.g. [26, 35])
$u_{n}^{(m+1)}$ $=$ $\left(\cos\beta u_{n+1}^{(m)}+i\sin\beta w_{n}^{(m)}\right)\exp(-\gamma_{n})\;\;\;\;\;\;$ (9)
$w_{n}^{(m+1)}$ $=$ $\left(i\sin\beta u_{n}^{(m)}+\cos\beta w_{n-1}^{(m)}\right)\exp(-\gamma_{n}).$ (10)
In the above equations, $u_{n}^{(m)}$ and $w_{n}^{(m)}$ are the pulse amplitudes at discrete time step $m$ and lattice site (unit cell) $n$ in the two fiber loops, $\beta$ is the coupling angle and $\gamma_{n}$ are uncorrelated stochastic losses (photonic traps) taking the two values $\gamma_{n}=\gamma>0$ with probability $p$ and $\gamma_{n}=0$ with probability $(1-p)$. A finite lattice, containing $L$ unit cells, is obtained by assuming a coupling angle $\beta=\pi/2$ at the lattice edges, so that in Eqs.(9) and (10) the site index $n$ can be restricted to the range $n=1,2,...,L$. Like for the continuous random walk with traps [Eq.(1)], the long-time dynamics of the discrete random walk defined by Eqs.(9) and (10) is governed by the eigenstate of the propagator $\mathcal{U}=\exp(-i\mathcal{H})$ with the longest lifetime, which is localized at the potential box of the Bernoulli distribution with the largest size $l=l_{m}$. An example of such an eigenstate, for a given realization of the trap distribution in the lattice, is shown in Fig.3(a). Such a state can be observed in the long-time random walk dynamics when one of the two fiber loops is seeded with a long flat-top optical pulse, corresponding to the initial condition $u_{n}^{(0)}=1/\sqrt{L}$ and $w_{n}^{(0)}=0$, as shown in Fig.3(b). In fact, the asymptotic temporal dynamics is dominated by the eigenstate of $\mathcal{U}$ with the longest lifetime, which is localized in the potential box of largest size $l_{m}$, as illustrated in Fig.3(a). Figure 3(c) shows the numerically-computed time evolution of the dynamical IPR, averaged over 200 realizations of disorder, clearly indicating stabilization to an asymptotic value which is in good agreement with the theoretical prediction [Eq.(8)].
Figure 3: (a) Numerically-computed behavior of the eigenstate (red circles) of the propagator $\mathcal{U}$ with the longest lifetime for the discrete-time photonic random walk with traps. The position of the traps in the lattice are denoted by the blue squares. Parameter values are $\beta=\pi/6$, $L=100$, $p=0.2$ and $\gamma=2$. (b) Temporal light dynamics in the synthetic lattice (behavior of $I_{n}=|u_{n}^{(m)}|^{2}+|w_{n}^{(m)}|^{2}$, normalized to its norm, versus propagation time step $m$ on a pseudocolor map) for uniform initial excitation $u_{n}^{(0)}=1/\sqrt{L}$, $w_{n}^{(0)}=0$. (c) Temporal evolution of the dynamical IPR versus time step $m$, averaged over 200 realizations of disorder. The IPR is defined as ${\rm IPR}=\sum_{n}I_{n}^{2}/\sum_{n}I_{n}$. The dashed horizontal curve is the asymptotic value predicted by Eq.(8).
Conclusion. We have demonstrated that Lifshitz tail states, extremely rare states emerging in the theory of Anderson localization, become significantly more prominent in non-Hermitian systems. These states play a crucial role in shaping the dynamical behavior of finite-size lattices, as illustrated through the study of photonic random walks with traps. Our findings highlight the importance of Lifshitz tails in non-Hermitian dynamics and provide a pathway to experimentally visualize extreme events, such as rare weakly localized states, in photonic platforms. This work opens new directions for exploring the impact of disorder and non-Hermiticity in various wave-based systems.
Disclosures. The author declares no conflicts of interest.
Data availability. No data were generated or analyzed in the presented research.
Funding. Agencia Estatal de Investigacion (MDM-2017-0711).
Supplemental document. See Supplement 1 for supporting content.